What are the standard time and date at (40°N, 110°W) when the standard time at (30°N, 125°E) is 1 A.M. on October 16?
Answer #1

East west separation = 110+125=235 degree

Time difference= 235*24/360= 15:40 hours

Standard at 110W is 9:20 AM , oct 15.